Assuming that radiation with a wavelength = 15.0 cm is used, that all the energy is converted to heat, and that 4.184 J is needed to raise the temperature of 1.00 g of water by 1.00 degree celcius, how many photons are necessary to raise the temperature of a 350 mL cup of water from 25 degree C to 95 degree C?
Answer #1
(photon energy)(number of photons) = (4.184 J/deg*g)*350g*70C
(h c/L)*N = 1.025*10^5 J
Solve for N.
L is the wavelength in meters
c = 3*10^8 m/s
h = Planck's constant
answered by: leya